OpenStudy (anonymous):

Using n = 4 equal-width rectangles, approximate the integral from negative 2 to 2 of the quantity x squared plus 8, dx. Use the left end-point of each sub-interval to determine the height of each rectangle.

jimthompson5910 (jim_thompson5910):

so you have this integral right? \[\Large \int_{-2}^{2}(x^2+8)dx\]

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

were you able to get the graph set up for this?

jimthompson5910 (jim_thompson5910):

if not, I can show you if you want?

OpenStudy (anonymous):

i dont know how

jimthompson5910 (jim_thompson5910):

ok the first step is to just graph f(x) all by itself. I'm using geogebra to do so

jimthompson5910 (jim_thompson5910):

we are given n = 4, a = -2 and b = 2 so \[\Large \Delta x = \frac{b-a}{n}\] \[\Large \Delta x = \frac{2 - (-2)}{4}\] \[\Large \Delta x = 1\] agreed?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

so what we do is start at a = -2 and count up by delta x units (ie count up by 1) until we hit b = 2

jimthompson5910 (jim_thompson5910):

so we have: -2, -1, 0, 1, 2 those will be x values to be plugged into the f(x) function

jimthompson5910 (jim_thompson5910):

for example, x = -2 f(x) = x^2 + 8 f(-2) = (-2)^2 + 8 f(-2) = 12 so the point (-2,12) is on the graph. Do the same for the other x values and you'll get these points

jimthompson5910 (jim_thompson5910):

the points A through D will determine the heights of the rectangles. The y coordinates of the points specifically

OpenStudy (anonymous):

(-2+8)+(-1+8)+(0+8)+(1+8)+(2+8)=40

jimthompson5910 (jim_thompson5910):

do you see in my last attached image how I set up the rectangles?

OpenStudy (anonymous):

yes i just saw it

jimthompson5910 (jim_thompson5910):

find the area of each rectangle, then add up the areas

OpenStudy (anonymous):

k

OpenStudy (anonymous):

12+9+8+9=38

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

Awesome than you so much.

jimthompson5910 (jim_thompson5910):

you're welcome